OpenStudy (anonymous):

can smeone give the general solution of y'= x+y/(2x+2y+3)

OpenStudy (anonymous):

it is under exact differential equations

OpenStudy (anonymous):

are u sure its not (x+y)/(2x+2y+3) ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

the take x+y=z and proceed. that is a useful substitution.

OpenStudy (anonymous):

ok will try

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

do u have an idea about exact differential equations

OpenStudy (anonymous):

yups but this substitution solves THIS particular sum. if you want a solution with exact differentials i can give u. but u may try this substitution.

OpenStudy (anonymous):

i have but the function will be in z and dy/dx then what should i do

OpenStudy (anonymous):

u will have \[\frac{dz}{dx}=\frac{3z+3}{2z+3}\] Integrate!

OpenStudy (anonymous):

dont understand as dy/dx=z/(2z+3) so did u get dz/dx

OpenStudy (anonymous):

use \[x+y=z \implies 1+\frac{dy}{dx}=\frac{dz}{dx}\] in case u r wondering how i got \[\frac{dz}{dx}\]

OpenStudy (anonymous):

oh u can do that, thanks, u made my day

OpenStudy (anonymous):

I solved it without using the concept of exact differential equations!

OpenStudy (anonymous):

yeah but is it ok to do that? will the answers be the same